# Recitation over Reasoning: How Cutting-Edge Language Models Can Fail on Elementary School-Level Reasoning Problems?

### v2

@@ Abstract
The rapid escalation from elementary school-level to frontier problems of the difficulty for LLM benchmarks in recent years have weaved a miracle for researchers that we are only inches away from surpassing human intelligence. However, is the LLMs’ remarkable reasoning ability indeed comes from true intelligence by human standards, or are they simply reciting solutions witnessed during training at an Internet level? To study this problem, we propose RoR-Bench, a novel, multi-modal benchmark for detecting LLM’s recitation behavior when asked simple reasoning problems but with conditions subtly shifted, and conduct empirical analysis on our benchmark. Surprisingly, we found existing cutting-edge LLMs unanimously exhibits extremely severe recitation behavior; by changing one phrase in the condition, top models such as OpenAI-o1 and DeepSeek-R1 can suffer $60\%$ performance loss on elementary school-level arithmetic and reasoning problems. Such findings are a wake-up call to the LLM community that compels us to re-evaluate the true intelligence level of cutting-edge LLMs.
@@
Figure 1: Panel a) shows an example of how current cutting-edge LLMs, OpenAI-o1-1217 [32], fails to address an elementary school-level math problem (see Appendix A.1 for the detailed response) with subtle but crucial condition change, simply reciting existing solution template (OpenAI-o1-1217 fails with input being either “floating” or “drifting”); panel b) shows the performance loss of cutting-edge LLMs due to reciting solution templates regardless of shifted conditions on our benchmark, which is a staggering $\sim 60\%$ score gap on simple reasoning and math problems.
@@ 1 Introduction
To better illustrate the existence of such cloud, here we examine a simple, GSM-8K [15] level math problem as an example in Fig. 1. Despite the simplicity of the problem, however, cutting-edge models such as OpenAI o1 [32] fails to solve such a problem; they simply recite the normal problem-solving paradigm of the problem, without carefully do the reasoning and check the subtle condition shift in the problem. With such phenomenon, we must ask the following tough question: Can the LLMs really solve simple reasoning problems, instead of simply reciting solution templates?
To find out the answer for this problem, in this work we propose RoR-Bench, a novel, multi-modal Chinese benchmark to detect the issue of Recitation over Reasoning for cutting-edge LLMs on simple reasoning problems, with $158$ pairs of text problems and $57$ pairs of image problems curated by humans; each pair consists of a simple, mostly elementary school-level reasoning problem and its variant with subtle but crucial condition shifts. We find that all cutting-edge LLM models have severe problem in reciting solutions instead of actually doing the reasoning, causing an accuracy loss that often exceeds $60\%$. Such phenomenon is particularly astounding on problems with no solutions; many cutting-edge LLMs, such as DeepSeek-R1, can even only recognize $<10\%$ cases as unsolvable. We explored initial solutions for mitigating the issue: adding notice prompts and providing subtly modified problems as few-shots. Although these solutions can mitigate the performance drop slightly, they are far from satisfactory and a more complete solution is still yet to be proposed.
@@ 3 RoR-Bench
In this section, we will introduce our proposed benchmark, RoR-Bench. RoR-Bench is a multimodal, question-answering Chinese benchmark consisting of pairs of problems, which are the original problems and the modified problems. The original problems are selected such that 1) cutting-edge LLMs can well-address, and 2) are mostly classic puzzles that appear in books and homework. The modified problems are created such that they look very similar to original problems, but with key condition modified and have completely different solution paradigms and answers. Fig. 2 provides an example for text and image problems in our benchmark.
Original problem: {CJK}UTF8gkai 两辆汽车从相距300公里的两个城市同时出发,相向而行。一辆车的速度是60公里/小时,另一辆车的速度是70公里/小时。请问它们需要多少小时才能相遇? (Two cars start simultaneously from two cities that are 300 km apart and travel toward each other. One car has a speed of 60 km/h, while the other has a speed of 70 km/h. How many hours will it take for them to meet?) Original answer: $300/(60+70)=\frac{30}{13}$.
Modified problem: {CJK}UTF8gkai 两辆汽车从相距300公里的两个城市同时出发,相而行。一辆车的速度是60公里/小时,另一辆车的速度是70公里/小时。请问它们需要多少小时才能相遇? (Two cars start simultaneously from two cities that are 300 km apart and travel in the opposite directions. One car has a speed of 60 km/h, while the other has a speed of 70 km/h. How many hours will it take for them to meet?) Modified answer: {CJK}UTF8gkai它们不会相遇。 (They will never meet.)
Refer to caption Original problem:{CJK}UTF8gkai这张图由多个同样的渐变菱形构成,它们整体看起来从上而下越来越暗,对吗? (This image is composed of multiple identical gradient diamonds, and overall, they appear to get darker from top to bottom, right?) Original answer:{CJK}UTF8gkai是的(马赫带效应) (Yes, it is a Mach band.) Modified problem: {CJK}UTF8gkai这张图由多个同样的渐变菱形构成,它们每个看起来从上而下越来越暗,对吗? (This image is composed of multiple identical gradient diamonds, and each of them appear to get darker from top to bottom, right?) Modified answer:{CJK}UTF8gkai不对,是自下而上 (No, it is from bottom to top.)
Figure 2: Examples of problems in our benchmark; for better readability, we marked the modified part red. Despite that we build a Chinese benchmark, OpenAI-o1-1217 [32] also fails with our English translation for these examples. See Appendix B.6 for more experiments on the English translation.
@@ 3.1 Dataset Curation
We asked 17 human annotators to collect simple reasoning problems from the Internet, mostly based on brain teaser collections in online blogs and sets of reasoning puzzles for children. Such problems become the original problems for our benchmarks. Then, we ask the annotators to modify the problems with the following instructions:
@@
    Different solution paradigm: The idea for addressing the modified problems must be completely different from the original problem. Simply changing numbers in the conditions (e.g. from 30km/h to 60km/h) is not allowed, as LLMs can well generalize to different figures in the condition.
@@
    No ambiguity: The modified problem must be rigorous, and only have one reasonable answer. For example, “how to cut a triangle cake into $4$ pieces (without any restrictions)” is too open to judge its correctness; “running competition in space (such that one cannot hear the starting gun)” is too ambiguous as humans cannot normally run in space, and LLMs may assume additional conditions such as the event is happening inside a space station.
@@
Each pair of original and modified problems will then be scrutinized by one of the $6$ moderators (or multiple moderators in borderline cases), to ensure that the problems have no error or duplication, and satisfy the principles above.
@@ 3.2 Dataset Statistics
RoR-Bench consists of a total of $215$ pairs of problems, with $158$ pairs of text problems and $57$ pairs of image problems. The image problems are all related to the property of the figure, while the text problem consists of $78$ math problems ($57$ arithmetic, $11$ geometry and $10$ probability / combinatorics) and $80$ reasoning problems ($38$ optimization, $10$ commonsense, $27$ deduction and $5$ game theory). See Fig. 3 for an illustration of the ratio for each type of problems. To ensure the simplicity of the problems, we curate the data such that all text inputs are less than $200$ characters, and each image problem only consists of a single image.
@@
Figure 3: An illustration of the types of the problem of our dataset.
@@ 4 Evaluations
In this section, we introduce the main results and empirical analysis for cutting-edge LLMs on our benchmark. In particular, we want to address the following questions: 1) Does the model really conduct reasoning over subtly modified conditions, or are they simply reciting existing solution paradigms to similar problems? If it is the latter, is it because the models view those changed conditions as typos (Sec. 4.1)? 2) Will simple fixes, such as using original problems as 1-shot, address the possible problem of recitation over reasoning (Sec. 4.2)? 3) How well does the LLMs perform when it comes to ill-posed problems, especially those with no solution (Sec. 4.3)?
4.1 Main Results
4.1.1 Text-based Problems
Evaluation. We evaluate 23 cutting-edge LLMs, which includes:
@@ 4.1.1 Text-based Problems
  • State-of-the-art small LLMs: Qwen-2.5-14B-Instruct [84] and Qwen-2.5-7B-Instruct.
As the answer to our question can be versatile with sometimes no solution, we do not adopt exact match as the metric. Instead, we use GPT-4o-1120 as the judge, which gives a binary (0/1) score (see Appendix A.2 for prompts) for LLM-generated answers. Each model is tested for $5$ times with temperature $0.7$ (we also report best-of-5 and greedy decoding results in Appendix B.1 and B.2 respectively). We use the average score (by GPT-4o-1120) as the metric over $5$ trials and $158$ problems, normalized to $0-100$; the higher score is the better.
@@
Table 1: Results on text-based problems of RoR-Bench, sorted by original score accuracy. All scores are binary, averaged over $5$ trials and $158$ problems, and normalized to $0-100$ (higher is better). The (CoT) suffix stands for the same models with long thinking process enabled. FC stands for “Forced Correct” prompt. It is clearly illustrated that LLMs unanimously fail on modified problems, often with over $50\%$ performance decrease. “Forced Correct” prompts somewhat helps, but is still far from addressing the recitation issue; also, the performance of original problems with “Forced Correct” prompts generally decrease, which indicates that adding prompt is not a valid solution.
@@
Results. Tab. 1 shows the result for all LLMs tested on RoR-Bench with original and modified problems, which shows a staggering $>50\%$ average performance decrease from scores on the original problems to the modified problems, and often $>60\%$ performance decrease for the best models such as DeepSeek-R1 and OpenAI-o3-mini-high. The best-of-5  [^3] Under the best-of-5 (Bo5) metric, the model is considered to get a score of $1$ if at least one of the $5$ trials get a score of $1$ under usual standards. With a low score but high Bo5, the model can be aligned with reinforcement learning [52] to quickly improve its score as positive samples are easy to acquire. performance of all LLMs also drop significantly (See Appendix B.1 for details), which indicates that such recitation issue is hard to be fixed simply by aligning techniques such as Reinforcement Learning (RL). Also, long thinking process does not seem to help for the modified problems. On modified problems, models such as DeepSeek-R1, OpenAI-o1-1217 and OpenAI-o3-mini-high works no better than those without long thinking process, such as GPT-4.5 Preview and Claude 3.7 Sonnet, despite having higher performance on original problems; also, Gemini-2.0 Flash-0121 and Claude 3.7 Sonnet works similarly on modified problems either with or without long thinking process. In spite of this, the performance on original problems and modified problems are positively related (Pearson correlation coefficient [53] $\rho\approx 0.72$), which indicates that the performance on modified problems are generally related to the base ability of the models.
@@
Is it because of the auto-correction of typos? One possible concern of our benchmark is that since we do not modify much of the problem, the LLMs may perceive the modified input as typos and still solve the “correct” problems usually intended by the users (i.e., original problems). To address such concern, we further test LLMs with the “Forced Correct” (FC) prompt added to the beginning of the problem:
@@
Forced Correct (FC) prompt: {CJK}UTF8gkai请回答下面的问题。题目保证没有问题,请严格按照题目字面所写的问题回答。 (“Please answer the following problem. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem.”)
@@
The results are also illustrated in Tab. 1. Surprisingly, even with the FC prompt, LLMs still show on average $>45\%$ performance decrease on modified problems, suggesting that the problem cannot be simply treated as an auto-correction of typos. Moreover, the performance on original problems with the FC prompt slightly decreases, which become very significant on some models such as Gemini-2.0 Pro-0205. Upon examining the output, we found that LLMs often become too strict on the expression of the problems; for example, when asked whether a game is fair, LLMs will question the definition of “fairness” and refuse to give a definitive answer (see Appendix B.5 for an example). Such result shows that simply adding prompts is not a valid way to address the recitation issue.
4.1.2 Vision-based Problems

### v3

@@ Abstract
The rapid escalation from elementary school-level to frontier problems of the difficulty for LLM benchmarks in recent years seems to bring us close enough to the “last exam” for LLMs to surpass humanity. However, is the LLMs’ remarkable reasoning ability indeed coming from true intelligence by human standards, or are they actually reciting solutions witnessed during training at an Internet level? To study this problem, we propose RoR-Bench, a novel, multi-modal benchmark for detecting LLM’s recitation behavior when asked simple reasoning problems but with conditions subtly shifted, and conduct empirical analysis on our benchmark. Surprisingly, we found existing cutting-edge LLMs unanimously exhibits extremely severe recitation behavior; by changing one phrase in the condition, top models such as OpenAI-o1 and DeepSeek-R1 can suffer $60\%$ performance loss on elementary school-level arithmetic and reasoning problems. Such findings are a wake-up call to the LLM community that compels us to reevaluate the true intelligence level of cutting-edge LLMs.
@@
Figure 1: Panel a) shows an example of how current cutting-edge LLMs, such as OpenAI-o1-1217 [33], OpenAI-o3 [51] and Gemini-Pro 2.5 [16] fails to address an elementary school-level math problem (see Appendix C.2 for links to the response) with subtle but crucial condition change, simply reciting existing solution template; panel b) shows the performance loss of cutting-edge LLMs due to reciting solution templates regardless of shifted conditions on our benchmark, which is a staggering $\sim 60\%$ score gap on simple reasoning and math problems.
@@ 1 Introduction
To better illustrate the existence of such cloud, here we examine a simple, GSM-8K [15] level math problem as an example in Fig. 1. Despite the simplicity of the problem, however, cutting-edge models such as OpenAI o1 [33] fails to solve such a problem; they simply recite the normal problem-solving paradigm of the problem, without carefully doing the reasoning and checking the subtle condition shift in the problem. With such phenomenon, we must ask the following tough question: Can the LLMs really solve simple reasoning problems, instead of simply reciting solution templates?
To find out the answer for this problem, in this work we propose RoR-Bench, a novel, multi-modal Chinese benchmark to detect the issue of Recitation over Reasoning for cutting-edge LLMs on simple reasoning problems, with $158$ pairs of text problems and $57$ pairs of image problems curated by humans; each pair consists of a simple, mostly elementary school-level reasoning problem and its variant with subtle but crucial condition shifts. We find that all cutting-edge LLM models have severe problem in reciting solutions instead of actually doing the reasoning, causing an accuracy loss that often exceeds $60\%$. Such phenomenon is particularly astounding on problems with no solutions; many cutting-edge LLMs, such as DeepSeek-R1, can even only recognize $10\%$ cases as unsolvable. We explored initial solutions for mitigating the issue: adding notice prompts and providing subtly modified problems as few-shots. Although these solutions can mitigate the performance drop slightly, they are far from satisfactory and a more complete solution is still yet to be proposed.
@@ 3 RoR-Bench
In this section, we will introduce our proposed benchmark, RoR-Bench. RoR-Bench is a multimodal, question-answering Chinese benchmark consisting of pairs of problems, which are the original problems and the modified problems. The original problems are chosen such that 1) cutting-edge LLMs can well-address, and 2) are mostly classic puzzles in books and homework. The modified problems are created such that they look very similar to original ones, but with key condition changed and have completely different solution paradigms and answers. Fig. 2 provides an example for text and image problems in RoR-Bench.
Original problem: 某警官发现前方100米处有一匪徒。警官赶紧以每秒5米的速度追。已知小偷的跑步速度为3米/秒,多少秒后警官可以追上这个匪徒? (A police officer spotted a thief 100 meters ahead of him. The officer started chasing the thief at 5 m/s. The thief runs at 3 m/s. How long does it take for the officer to catch the thief?) Original answer: 100/(53)=50100/(5-3)=50s.
Modified problem: 某警官发现前方100米处有一匪徒,匪徒没有发现警官。警官赶紧以每秒5米的速度追,已知小偷的跑步速度为3米/秒,多少秒后警官可以追上这个匪徒? (A police officer spotted a thief 100 meters ahead of him, but the thief did not notice the officer. The officer started chasing the thief at 5 m/s. The thief can run at 3 m/s. How long does it take for the officer to catch the thief?) Modified answer: $100/5=20$s.
Refer to caption Original problem:这张图由多个同样的渐变菱形构成,它们整体看起来从上而下越来越暗,对吗? (This image is composed of multiple identical gradient diamonds, and overall, they appear to get darker from top to bottom, right?) Original answer:是的(马赫带效应) (Yes, it is a Mach band.) Modified problem: 这张图由多个同样的渐变菱形构成,它们每个看起来从上而下越来越暗,对吗? (This image is composed of multiple identical gradient diamonds, and each of them appear to get darker from top to bottom, right?) Modified answer:不对,是自下而上 (No, it is from bottom to top.)
Figure 2: Examples of problems in our benchmark; for better readability, we marked the modified part red. Despite that we build a Chinese benchmark, OpenAI-o1-1217 [33], OpenAI-o3 [51] and Gemini 2.5 Pro [16] all fail with our English translation for these examples. See Appendix B.2 for another example and Appendix C.3 for links to experiment records on the English translation.
@@ 3.1 Dataset Curation
We asked 17 human annotators (all native speakers to ensure dataset quality) to collect simple reasoning problems from the Internet, mostly based on brain teaser collections in online blogs and sets of reasoning puzzles for children. Such problems become the original problems for our benchmarks. Then, we ask the annotators to modify the problems with the following instructions:
@@
    Different solution paradigm: The idea for addressing the modified problems must be completely different from the original problem. Simply changing numbers in the conditions (e.g. from 30km/h to 60km/h) is not allowed, as LLMs can well generalize to different figures in the condition. The modified problem is often simpler and more straightforward; for example, the modification can be “how to discriminate the two items in a black box” to “how to discriminate the two items in a transparent box”.
@@
    No ambiguity: The modified problem must be rigorous, and only have one reasonable answer. For example, “how to cut a triangle cake into $4$ pieces (without any restrictions)” is too open to judge its correctness; “running competition in space (such that one cannot hear the starting gun)” is too ambiguous as humans cannot normally run in space, and LLMs may assume additional conditions such as the event is happening inside a space station. Note, both the collected original problems and the modified problems are intended to be easy to solve, with the latter having unconventional conditions.
@@
Each pair of original and modified problems will then be scrutinized by one of the $6$ moderators (or multiple moderators in borderline cases), to ensure that the problems have no error or duplication, do not contain any identifying or offensive content, and satisfy the principles above.
@@ 3.2 Dataset Statistics
RoR-Bench consists of a total of $215$ pairs of problems, with $158$ pairs of text problems and $57$ pairs of image problems. Such size is comparable to the most related works, e.g. MATH-Perturb [26] with $279$ pairs of problems, and MathTrap [95] with $105$ original public triplet of problems [^2] The rest 895 are paraphrases by GPT..
The image problems are all related to the property of the figure, while the text problem consists of $78$ math problems ($57$ arithmetic, $11$ geometry and $10$ probability / combinatorics) and $80$ reasoning problems ($38$ optimization, $10$ commonsense, $27$ deduction and $5$ game theory). See Fig. 3 for an illustration of the ratio for each type of problems. To ensure the simplicity of the problems, we curate the data such that all text inputs are less than $200$ characters, and each image problem only consists of a single image.
@@
Figure 3: An illustration of the types of the problem of our dataset, which covers a variety of reasoning problems; we double-checked the problems to ensure the low difficulty of the original ones.
@@ 4 Evaluations
In this section, we introduce the main results and empirical analysis for cutting-edge LLMs on RoR-Bench. In particular, we aim to address the following questions: 1) Does the model really conduct reasoning over subtly modified conditions, or are they simply reciting existing solution paradigms to similar problems? If it is the latter, is it because the models view those changed conditions as typos (Sec. 4.1, Sec. 4.2)? 2) Will simple fixes, such as using original problems as 1-shot, address the possible problem of recitation over reasoning (Sec. 4.3)? 3) How well can the LLMs deal with ill-posed problems, especially those with no solution (Sec. 4.4)? 4) In general, why does the recitation phenomenon happen (Sec. 4.5)?
4.1 Text-based Problems
Evaluation. We evaluate 23 cutting-edge LLMs, which include:
@@ 4.1 Text-based Problems
  • Small LLMs: Qwen-2.5-14B-Instruct [85] and Qwen-2.5-7B-Instruct.
As the answer to our question can be versatile with sometimes no solution, we do not adopt exact match as the metric. Instead, we use GPT-4o-1120 as the judge, which gives a binary (0/1) score (see Appendix B.1 for prompts) for LLM-generated answers. Each model is tested for $5$ times with temperature $0.7$, following default by OpenAI API reference document [52]; we choose non-greedy decoding to test more rollouts and better differentiates the models’ perormance. We also report best-of-5 and greedy decoding results in Appendix C.4 and C.5 respectively). We use the average score (by GPT-4o-1120) as the metric over $5$ trials and $158$ problems, normalized to 01000-100; the higher score is the better.
@@
Table 1: Results on text-based problems of RoR-Bench, sorted by original score accuracy. All scores are binary, averaged over $5$ trials and $158$ problems, and normalized to 01000-100 (higher is better). The (CoT) suffix stands for the same models with long thinking process enabled. FC stands for “Forced Correct” prompt. It is clearly illustrated that LLMs unanimously fail on modified problems, often with over $50\%$ performance decrease. “Forced Correct” prompts somewhat helps, but is still far from addressing the recitation issue; also, the performance of original problems with “Forced Correct” prompts generally decrease, which indicates that adding prompt is not a valid solution.
@@
Results. Tab. 1 shows the result for all LLMs tested on RoR-Bench with original and modified problems, which shows a staggering $50\%$ average performance decrease from scores on the original problems to the modified problems, and often $60\%$ performance decrease for the best models such as DeepSeek-R1 and OpenAI-o3-mini-high. The best-of-5  [^4] Under the best-of-5 (Bo5) metric, the model is considered to get a score of $1$ if at least one of the $5$ trials get a score of $1$ under usual standards. With a low score but high Bo5, the model can be aligned with reinforcement learning [53] to quickly improve its score as positive samples are easy to acquire. performance of all LLMs also drop significantly (See Appendix C.4 for details), which indicates that such recitation issue is hard to be fixed simply by aligning techniques such as Reinforcement Learning (RL). Also, long thinking process does not seem to achieve better performance. On modified problems, models such as DeepSeek-R1, OpenAI-o1-1217 and OpenAI-o3-mini-high works no better than those without long thinking process, such as GPT-4.5 Preview and Claude 3.7 Sonnet, despite having higher performance on original problems; also, Gemini-2.0 Flash-0121 and Claude 3.7 Sonnet works similarly on modified problems either with or without long thinking process. In spite of this, the performance on original problems and modified problems are positively related (Pearson correlation coefficient [54] $\rho\approx 0.72$), which indicates that the performance on modified problems are generally related to the base ability of the models.
4.1.1 Reliability analysis of the LLM judge
While using LLM judges can potentially introduce bias, we select LLM judges for two reasons:
  • A large portion of problems in our benchmark (especially the non-math ones) are inherently very hard to be verified by rules (e.g., “How can A ensure victory in a game?” or “How can we quickly find some object in a set of objects?”).
  • The judge also considers whether the LLM recognizes the trap in the modified questions, which is important on deciding whether the model is reciting solution paradigms. As most of our vision-based problems are binary multiple choices, verifiable answers could lead to higher scores with guessing.
We manually verified the effectiveness of the current LLM judge on OpenAI-o1-1217. We did not find any incorrect judging result for vision-based problems, as we either use exactly the same problem but modify the figure or ask for different properties of the same figure. For text-based problems, we find 20 out of 157*2=314 (original vs modified) arguable cases, where in 12 cases o1 somewhat realizes the subtle difference but made otherwise assumptions (e.g. “the problem may suggest xxx, but we should assume xxx”) [^5] We view such result as incorrect as “partially correct” confuses the LLM judge in practice., in 7 cases LLM gives a different valid answer (some of them are suboptimal as o1 is unaware of the problem modification), and in 1 case is the judge made a mistake itself. The judgment accuracy on modified text-based problems is no less than 90%, which validates the existence of performance gap between original and modified problems. We further test our models on a verifiable subset of our benchmark; see Appendix C.1 for results.
4.1.2 Excluding Auto-Correction of Typos
@@
One possible concern of our benchmark is that since we do not modify much of the problem, the LLMs may perceive the modified input as typos and still solve the “correct” problems usually intended by the users (i.e., original problems). To address such concern, we further test LLMs with the “Forced Correct” (FC) prompt added to the beginning of the problem:
@@
Forced Correct (FC) prompt: 请回答下面的问题。题目保证没有问题,请严格按照题目字面所写的问题回答。 (“Please answer the following problem. The problems are guaranteed to be correct; please strictly follow the literal meaning of the problem.”)
@@
The results are also illustrated in Tab. 1. Surprisingly, even with the FC prompt, LLMs still show on average $45\%$ performance decrease on modified problems, suggesting that the problem cannot be simply treated as an auto-correction of typos. Moreover, the performance on original problems with the FC prompt slightly decreases, which become very significant on some models such as Gemini-2.0 Pro-0205. Upon examining the output, we found that LLMs often become too strict and overthink on the expression of the problems; for example, when asked whether a game is fair, LLMs will question the definition of “fairness” and refuse to give a definitive answer (see Appendix C.8). Such result shows that simply adding prompts is not a valid solution to the recitation issue.
4.2 Vision-based Problems